OpenStudy (anonymous):

2. Suppose that vectors v1 = (1, 2), v2 = (2,−1), and that the basis B is B=v1, v2 . (this is a list) Let T be the linear transformation from R2 to R2 given by T(v1) = v1 and T(v2) = 0. (a) Write down the matrix for T in the new basis B. (You should be able to do this directly from the definition of T). (b) Use this to write down the matrix for T in the standard basis.

OpenStudy (amistre64):

1 2 2 -1 is B T(x) = Ax a c b d = A 1(a,b) + 2(c,d) = (1,2) 1a + 2c = 1 1b + 2d = 2 2(a,b)-1(c,d) = (0,0) 2a -1c = 0 2b - 1d = 0 Of course this inst according to the definition of a linear transform, but its how i do it

OpenStudy (amistre64):

1a + 2c = 1 2a - 1c = 0 *2 1a + 2c = 1 4a - 2c = 0 ---------- 5a = 1 ; a = 1/5 1a + 2c = 1 *-2 2a - 1c = 0 -2a -4c = -2 2a - 1c = 0 ----------- -5c = -2; c = 2/5

OpenStudy (amistre64):

1b + 2d = 2 2b - 1d = 0 -2b -4d = -4 2b - 1d = 0 ------------- -5d = -4 ; d=4/5 1b + 2d = 2 4b - 2d = 0 ------------- 5b = 2; b = 2/5 \[A=\frac{1}{5}\begin{pmatrix}1&2\\2&4\end{pmatrix}\] maybe might have to leave the 1/5 in there to work tho

OpenStudy (amistre64):

B = v1, v2 newB = T(v1), T(v2) \[T\binom{1}{2}+ T\binom{2}{-1}\] \[1\binom{1}{2}+2\binom{1}{-1/2}\] ugh ... i just dont have the practice at doing it htis way :/

OpenStudy (anonymous):

ok so i'm confused, the matrix that I got for a was [1 0] [0 0]

OpenStudy (amistre64):

howd you get it? that might help me know if im completely off

OpenStudy (anonymous):

I got it because if T(v1) is (1, 2) and T(v2) is (0, 0), then the matrix that takes the vectors to that one is [1 0] [0 0]

OpenStudy (anonymous):

but i could be wrong, which is exactly why i posted the question here haha.

OpenStudy (amistre64):

1. 1 2. 0 = 1+0 1 0 0 = 0+0 not equal 2 so the transformation matrix doesnt tramsform v1 into v1

OpenStudy (anonymous):

lmao ok then i'm wrong.

OpenStudy (anonymous):

but that was my mind set for trying to get the right answer..

OpenStudy (amistre64):

\[A=\frac{1}{5}\begin{pmatrix}1&2\\2&4\end{pmatrix}\binom{1}{2}=1\binom{1}{2}+2\binom{2}{4}=\binom{1+4=5}{2+8=10}/5=\binom{1}{2}\] right?

OpenStudy (anonymous):

right

OpenStudy (amistre64):

lets see if it works for v2 :) \[A\vec{v_2}=\frac{1}{5}\begin{pmatrix}1&2\\2&4\end{pmatrix}\binom{2}{-1}=2\binom{1}{2}-1\binom{2}{4}=\binom{2-2=0}{4-4=0}/5=\binom{0}{0}\]

OpenStudy (amistre64):

how to get to that thru just using the definitions of linear transformations I cant tell ...

OpenStudy (amistre64):

i take a generic A matrix, and create 2 sets of equations to solve

OpenStudy (anonymous):

ok, now do you know how i would find the stand basis matrix for that?

OpenStudy (amistre64):

i think that is it, but im reading up on it at the moment to be sure

OpenStudy (amistre64):

the part "a" is spose to get you to the A matrix that I did; but thru a different technique i think

OpenStudy (anonymous):

hmm, oh. i dont think so, part a asks for the matrix in the basis B

OpenStudy (amistre64):

B = 1 2 2 -1 newB = 1 0 2 0 is that part a?

OpenStudy (amistre64):

if i could see your textbook im sure id be like "doh!!"

OpenStudy (amistre64):

1.1 +0. 0 = 1 2 0 0 0.1 +1. 0 = 0 2 0 0 which is why you came up with the : 1 0 0 0 to begin with

OpenStudy (anonymous):

wahhh i wish i knew. that makes sense.

OpenStudy (amistre64):

ill read up on it tonight and see what I can remember :) good luck

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

did you ever end up figuring anymore of it out? i'm so stumped.

OpenStudy (amistre64):

nothing really that makes any sense to me .... B = v1 v2 B = 1 2 2 -1 .......................................... newB = T(v1) T(v2) newB = 1 0 2 0 seems to be the way its looking from what I can gather. that should be part a its the part b thats got me frazzled

OpenStudy (anonymous):

ok so all of the stuff you put yesterday with everything over 5, that matrix, that has nothing to do with it? haha

OpenStudy (amistre64):

from what ive gathered; it was not the manner in which they are wanting this accomplished

OpenStudy (amistre64):

im not sure what newB and B have to do with each other rather than writing them up as baseses for vector spaces

OpenStudy (amistre64):

what is the name of the chapter that you are working on?

OpenStudy (anonymous):

it's all about changing basis... idk if you gathered that. so basically its just asking for the matrix of the transformation with respect to basis b and then in part b its asking for that same matrix only with respect to the standard basis. they're homework questions assigned and not in the textbook.

OpenStudy (amistre64):

so, part b is what: T(v1) = (1,0) ; T(v2) = (0,1)?

OpenStudy (anonymous):

i'm not sure if that's how you write it ... because i dont know if it's neccessarily going to the standard matrix, but what would the standard basis matrix be of the transformation .. if that makes any sense.

OpenStudy (amistre64):

i aint got a clue yet :) B = v1 v2 \[B\vec{x} =x_1v_1+x_2v_2\] \[B\vec{x} =x_1\binom 12+x_2\binom 2{-1}\] \[B' = T(\vec{v_1})\ T(\vec{v_2})\] \[B'\ \vec{x} =x_1T(v_1)+x_2T(v_2)\] \[B'\ \vec{x} =x_1\binom 12+x_2\binom 00\] what this has to do with anything I cant tell

OpenStudy (anonymous):

lmao its all good, it's like the blind leading the blind here! thanks anyways!

OpenStudy (amistre64):

yeah, all in all, good luck with it ;)